OpenStudy (anonymous):

log square root p/r^2 how can you put it in an equation using only log p and/or log r???

OpenStudy (sandra):

hey check out this guide: http://abacus.bates.edu/acad/acad_support/msw/exps_and_logs.pdf

OpenStudy (sandra):

also, can you try using parens (or the equation editor in the lower left) to write it out? a bit hard to read in your top line

OpenStudy (sandra):

\[\log_{\sqrt{p/r ^{2}}} \] ?

OpenStudy (sandra):

\[\log \sqrt{p/r ^{2}}\] or this

OpenStudy (anonymous):

\[\log (\sqrt{P} \div R ^{2})\]